OpenStudy (anonymous):

give the value of 3 in this problem 543,600,211

jimthompson5910 (jim_thompson5910):

The value 3 is in the millions place, so 3 really represents 3,000,000 (ie 3 million)

OpenStudy (anonymous):

whats 129 divide 9

jimthompson5910 (jim_thompson5910):

You can divide both by 3 to get \[\frac{129}{9}=\frac{43}{3}\] Then use long division (or a calculator) to get \[\frac{43}{3}=14.\bar{3}\] So\[\frac{129}{9}=14.\bar{3}\]

OpenStudy (anonymous):

thanks